OpenStudy (anonymous):

What is the interest on $950 invested for 4 years in an account that earns simple interest at a rate of 3% per year? Can someone answer and explain?

OpenStudy (anonymous):

every year the interest added is 3% of $950 = ( 3* 950) / 100 = $28.5 so total interest added in 4 years is 4 * 28.50 = $114